My friend builds a fish pond for his fishes, goldfishes and a variety of carp. He decided to use two different hoses to easily fill his fish pond with water. If he used the thin hose, he would able to fill the pond in 24 minutes. But if he used the thick one, able to fill the pond in 8 minutes. Using the two hoses together, how long will it take him to fill the fishpond?
Found 2 solutions by Solver92311, greenestamps:
Answer by Solver92311(821)



If entity A can complete an entire job in time periods, then entity A can complete of the job in one time period. Likewise, entity B can complete of the job in one time period if B can complete the whole job in time periods. So if is the time it takes both working together to complete the job, is the fraction of the job both working together can complete in one time period. In short:



So plug in the values you have for and and solve for

John

My calculator said it, I believe it, that settles it

From
I > Ø

Answer by greenestamps(13200)



Here is an easy alternative to the standard algebraic method shown by the other tutor.

(1) Determine the least common multiple of the two given times. In this example, the LCM of 8 minutes and 24 minutes is 24 minutes.

(2) Calculate the number of times each hose could fill the pond in 24 minutes. The thin hose could fill it 24/24=1 time in 24 minutes; the other hose could fill it 24/8=3 times in 24 minutes.

(3) Together the two hoses could fill the pond 1+3=4 times in 24 minutes, so the two together could fill the pond once in 24/4=6 minutes.

ANSWER: 6 minutes